### v1

# Improving LLM-based Global Optimization with Search Space Partitioning

Abstract
Large Language Models (LLMs) have recently emerged as effective surrogate models and candidate generators within global optimization frameworks for expensive blackbox functions. Despite promising results, LLM-based methods often struggle in high-dimensional search spaces or when lacking domain-specific priors, leading to sparse or uninformative suggestions. To overcome these limitations, we propose HOLLM, a novel global optimization algorithm that enhances LLM-driven sampling by partitioning the search space into promising subregions. Each subregion acts as a “meta-arm” selected via a bandit-inspired scoring mechanism that effectively balances exploration and exploitation. Within each selected subregion, an LLM then proposes high-quality candidate points, without any explicit domain knowledge. Empirical evaluation on standard optimization benchmarks shows that HOLLM consistently matches or surpasses leading Bayesian optimization and trust-region methods, while substantially outperforming global LLM-based sampling strategies.
1 Introduction and Motivation
Global optimization (jones98, ; rios2013, ) (also known as gradient-free or zeroth-order optimization) of blackbox functions, where the only information provided to the optimizer is the function value, is a fundamental challenge across numerous domains including hyperparameter tuning (snoek12, ; turner21a, ), policy search (calandra16, ), molecular design and chemical engineering (langer04, ; lobato17, ), just to name a few. Methods such as Bayesian optimization (Shahriari2016, ; garnett_bayesoptbook_2023, ) and evolutionary algorithms (hansen2016evolution, ) have been a standard and effective choice across various applications. However, they typically require assumptions regarding the underlying objective function’s nature, which consecutively affect algorithmic design choices.
At the same time, recent advances in Large Language Models (LLMs) have demonstrated remarkable capabilities in generative modelling and reasoning (brown20, ; gpt4, ; touvron2023, ), suggesting their potential usage for optimization tasks as well (song24, ). Efforts in integrating LLMs within blackbox optimization algorithms as surrogate models or as candidate samplers have already shown encouraging results (yang2024large, ; liu2024large, ; zhang2023using, ; aglietti2025funbo, ; agarwal2025searching, ; kristiadi24, ). Yet these methods typically rely on carefully engineered, domain-specific prompts, and in higher dimensions and complex search spaces the LLM’s suggestions tend to scatter sparsely, covering only a fraction of the domain (krishnamurthy24, ).
As a motivating example, we investigated the capabilities of LLMs to simulate uniform sampling from a unit hypercube. In Figure 1(a) we show 80 samples drawn from the unit square $[0,1]^{2}$, comparing uniform sampling (blue) with Gemini-1.5’s (gemini15, ) attempt at simulating uniform sampling using the prompt provided in Listing 1 in Appendix D (green points), and Gemini-1.5 performing uniform sampling with 5 samples per smaller subregion, using the same prompt (red points). We can clearly notice that even in 2D the LLM demonstrates high bias when sampling, therefore failing to appropriately fill the space as it was tasked to, whilst partitioning the space and prompting the LLM 16 times yields a more faithful simulation. Another illustrative example is shown in Figure 1(b), where we prompt Gemini-1.5 (using the prompt shown in 2) to sample close to the two global minima (red stars) of a quadratic function, given the input space boundaries. We can clearly notice the higher sampling bias when the input space is $[0,1]^{2}$ instead of the smaller regions denoted via the dashed bounding boxes. Finally, in Figure 1(c), we compute the Hausdorff distance, $d_{H}(\mathcal{P},[0,1]^{8})$, between the set of $N\in\{10,20,50,70,100\}$ sampled points $\mathcal{P}$ and the 8-dimensional unit hypercube $[0,1]^{8}$. The blue curve indicates the values for standard uniform sampling and the other ones performed by various non-agentic LLMs. Similarly as in the 2D case, partitioning the hypercube into 32 regions and sampling within each (Gemini 1.5 + partitioning) notably improves the spatial coverage, enabling the LLM to more closely approximate uniform sampling.
Figure 1: (a) 80 samples in $[0,1]^{2}$: Gemini-1.5 simulating uniform sampling (green), and with region-wise partitioning (red) using the prompt in Listing 1. (b) Gemini-1.5 prompted (see Listing 2) to generate 80 samples around the 2 minima (red crosses) globally (triangles) and withing the two bounding boxes (circles). (c) Hausdorff distance $d_{H}(\mathcal{P},[0,1]^{8})$ for uniform vs. LLM-simulated sampling in the 8-D hypercube.
In this paper, we introduce Hierarchical Optimization with Large Language Models (HOLLM), a novel blackbox optimization method that leverages adaptive spatial partitioning to guide LLM-based sampling. HOLLM iteratively builds a KD-tree on existing evaluation data, creating adaptive local partitions whose granularity evolves with sampling density. Each subregion is assigned a bandit-inspired utility score, balancing exploitation (regions with promising observed values) and exploration (geometrically large or statistically uncertain regions). Subregions are selected stochastically according to these scores, and LLMs then generate localized candidate proposals within the chosen regions. As LLMs trained on optimization literature and scientific data encode a valuable meta-prior about typical function behavior (e.g., local unimodality), we effectively harness this prior without assuming a fixed parametric surrogate (e.g., Gaussian Process). Furthermore, restricting candidate generation to smaller, lower-dimensional subregions significantly reduces LLM sampling difficulty compared to global high-dimensional sampling. We note that spatial partitioning heuristics have already proven effective in continuum-armed bandits (munos11, ; bubeck11, ; valko13, ; grill15, ), Trust Region Bayesian optimization (eriksson19, ; daulton22, ), and Monte Carlo Tree Search (kim20, ; wang20, ; yang21, ). The key contribution of this work is the integration of these partitioning ideas to substantially improve LLM-driven global optimization performance.
Empirical evaluations on continuous and discrete benchmark functions, including hyperparameter optimization and neural architecture search, demonstrate that HOLLM effectively balances exploration and exploitation, matching or outperforming state-of-the-art methods, including established Bayesian optimization variants and Trust Region algorithms, particularly in scenarios requiring efficient navigation of complex landscapes. Furthermore, compared to approaches that prompt the LLM to propose candidates globally, HOLLM achieves considerable gains by focusing LLM suggestions locally. We provide the implementation of our algorithm in the following repository: https://github.com/automl/hollm.
2 Background and Related Work
We consider the problem of maximizing a blackbox function $f:\mathcal{X}\rightarrow\mathbb{R}$ where $\mathcal{X}$ is a compact domain. The objective is to find $x^{*}=\operatorname*{arg\,max}_{x\in\mathcal{X}}f(x)$ through a sequence of function evaluations. In this blackbox setting, we do not have access to gradients or other properties of $f$, and can only observe function values at queried points. The performance of optimization algorithms in this context can also be measured using simple regret or cumulative regret. For a sequence of evaluated points $x_{1},x_{2},\ldots,x_{t}$, the simple regret after $t$ iterations is defined as: $r_{t}=f(x^{*})-\max_{i\in\{1,\ldots,t\}}f(x_{i})$, while the cumulative regret is: $R_{t}=\sum_{i=1}^{t}(f(x^{*})-f(x_{i}))$.
Bayesian Optimization. Bayesian Optimization (BO) (garnett_bayesoptbook_2023, ; Frazier2018ATO, ; Shahriari2016, ) is a well-established framework for optimizing expensive blackbox functions by maintaining a probabilistic surrogate (typically a Gaussian Process (Rasmussen2006Gaussian, )) to guide evaluations and optimizing an acquisition function (e.g. Expected Improvement (zhan20, )) in order to balance exploration and exploitation and efficiently search the space. Extensions like TuRBO (eriksson19, ; daulton22, ) address high-dimensional settings by maintaining multiple trust regions, which are dynamically resized based on optimization progress, enabling scalable and focused exploration around promising evaluations via local GPs.
Multi-Armed Bandits and Hierarchical Optimization Algorithms. Multi-Armed Bandits (MABs) (mab-suvey, ) deal with the problem of sequential decision-making under the exploration-exploitation dilemma. In the basic setting, a MAB algorithm repeatedly selects among a fixed (also infinite) number of arms or actions, each with an unknown reward distribution, aiming to minimize the cumulative regret. In the global optimization setting, the arms are the points that lie in the input space $\mathcal{X}$ and at each iteration $t$, an arm $x_{t}\in\mathcal{X}$ is pulled and the regret is computed by evaluating the function $f(x_{t})$ (grill15, ). Several MAB algorithms leverage hierarchical space partitioning (kleinberg08, ). Most notably, HOO (bubeck11, ) constructs a hierarchical partitioning of the search space using $n$-ary trees and at each step, an unexplored region (tree leaves) is selected based on upper confidence bounds (UCB) (auer02, ; Kocsis2006, ) and $f$ is evaluated at a point uniformly sampled inside the selected region. Building on HOO, extensions include parallel versions (grill15, ), optimization without explicit smoothness knowledge (munos11, ) or under noisy observations (valko13, ), and adaptive trees (bull13, ) including Monte Carlo bandits (wang20, ; yang21, ). Most of these methods come with theoretical guarantees on regret bounds that depend on the dimensionality and smoothness properties of the objective function.
Large Language Models for Blackbox Optimization. Recent work has increasingly explored integrating LLMs into blackbox optimization workflows. Some approaches prompt LLMs directly to generate candidate solutions in natural language (liu2024large, ; yang2024large, ; agarwal2025searching, ; zhang2023using, ), use them to estimate uncertainty (ramos23, ), extract features (kristiadi24, ), or even design novel acquisition functions (aglietti2025funbo, ). Others replace traditional surrogate models with LLMs to predict function values and guide search in applications such as hyperparameter tuning (liu2024large, ) and molecular design (ramos23, ). However, these methods often rely on carefully engineered prompts containing domain-specific information (e.g., dataset statistics or problem descriptions), raising concerns about their robustness in domains where this information is not available. Recent work by krishnamurthy24 shows that, in simple MAB settings, LLMs struggle to explore effectively without significant prompt intervention, highlighting their limitations in decision-making.
Our algorithm builds upon these foundations in order to improve LLM-based blackbox optimization by integrating tree-based space partitioning, a UCB-inspired score function for balancing exploration and exploitation, and LLM-based candidate generation within locally promising regions.
3 HOLLM: Hierarchical Optimization with LLMs
In this section, we present the HOLLM algorithm for optimizing potentially noisy blackbox functions $f:\mathcal{X}\rightarrow\mathbb{R}$, which consists of 5 main steps: Partition, Score, Select, Sample and Evaluate. Given an initial set of $n_{0}$ evaluations $\mathcal{D}_{n_{0}}=\{(x_{i},f(x_{i}))\}_{i=1}^{n_{0}}$, the algorithm iteratively calls each of these steps. It starts by adaptively partitioning the search space in a data-driven way, scores each of these regions to balance exploration-exploitation, selects the $M$ most promising regions based on their score, leverages LLMs to sample candidates within these regions, and finally evaluates the best candidates according to their predicted function value from the LLM. We provide an illustrative depiction of these steps in Figure 2. This approach allows the LLM to focus on promising smaller regions of the space while benefiting from the global partitioning strategy. We provide the algorithm pseudocode in Algorithm 1 and a more detailed version in the Appendix A. In the following, we explain each step in detail.
3.1 Partition: Adaptive Discretization
Based on the motivating examples we presented in Section 1, we hypothesize that firstly identifying promising smaller regions in the input space $\mathcal{X}$ makes the LLM-based sampling more reliable compared to prompting the LLM to sample globally. To this end, we propose using an adaptive input space partitioning method based on the evaluated data at each iteration of the algorithm. In order to obtain disjoint space partitions that cover the entire space, we use k-dimensional trees (KD-trees), a space-partitioning data structure that recursively divides the space into half-spaces, so that we can efficiently compute the partitions in high dimensions ($\mathcal{O}(t\log(t))$ for a balanced tree where $t$ is the number of iterations), whereas for other methods, such as a Delaunay triangulation (gramacy22, ) and Voronoi diagram (kim20, ; Wycoff2024VoronoiCF, ), this would become quickly impractical as the dimension $d$ increases. Each non-leaf node in a KD-tree represents a splitting hyperplane perpendicular to one of the coordinate axes, dividing the space into two parts. Points on the "left" side of this hyperplane are represented by the left subtree, and points on the "right" side are represented by the right subtree. Starting from the root node $\mathcal{X}_{\varnothing}\!=\!\mathcal{X}$, every internal node chooses a split dimension $s$ (the one with the largest variance among points in the node) and a split value $\delta$ (the mean across the selected dimension). This produces two child nodes
$\mathcal{X}_{\text{left}}=\{x\in\mathcal{X}:x_{s}\leq\delta\},\quad\mathcal{X}_{\text{right}}=\{x\in\mathcal{X}:x_{s}>\delta\},$
whose union equals their parent and whose interiors are disjoint. After inserting $n$ sample points, the $K$ leaves $\{\mathcal{X}_{l}\}_{l=1}^{K}$ form a partition of $\mathcal{X}$ into axis-aligned hyperrectangles and contain information about the points evaluated within it, including their coordinates and function values. We denote the set of indices each leaf $\mathcal{X}_{\ell}$ holds as: $I_{\ell}=\{\,i\leq t:x_{i}\in\mathcal{X}_{\ell}\}$, with sample size $n_{\ell}=|I_{\ell}|\leq m_{t}$. $m_{t}$ is the maximum number of points a leaf in the KD-tree can keep before splitting, parameterized by the number of iterations. At the start of round $t$, we optionally set $m_{t}=m_{0}+\bigl{\lceil}\lambda\log(1+t)\bigr{\rceil}$, where $m_{0}$ ($\lceil d/2\rceil$ by default) is the initial leaf size and $\lambda$ (00 by default) is the growth parameter. The logarithmic growth of $m_{t}$ ensures that the partitions do not become too fine-grained quickly.
An infinite-armed bandit view. Conceptually, the KD-tree can be interpreted as a data-driven discretizer in infinite-armed bandits (wang08, ; carpentier15, ): its leaves form a coarse partition at the beginning of learning and refine only where information accumulates, mirroring the “zooming” phenomenon in continuum–armed bandits (kleinberg08, ; wang08, ; munos11, ; bubeck11, ; slivkins11, ; valko13, ; bull13, ; carpentier15, ; grill15, ). A key distinction, however, is that the entire tree is re-fitted at every round, adapting the partitions’ boundaries based on the current data to avoid potential early convergence to local minima. This strategy is similar to wang20 ; yang21 , where they observed that recomputing the partitioning every few iterations resulted in better empirical performance. Although partition boundaries may merge or shift, one can frame the procedure as operating on a fixed, infinite KD-tree whose internal nodes are activated and deactivated on the fly, as abstracted in adaptive-treed bandits (bull13, ; slivkins11, ). Following this paradigm, every axis-aligned hyper-rectangular partition, $\mathcal{X}_{l}\subset\mathcal{X}$, can be seen as a "meta-arm" (slivkins11, ). The reservoir of all such boxes is uncountable, hence discovering arms, rather than merely pulling them, becomes part of the learning problem. In this language, our algorithm may be viewed as an infinite-armed bandit strategy that (i) repeatedly draws a batch of candidate active input space partitions by re-fitting a KD-tree to the ever-growing set of evaluations, and (ii) allocates pulls among those boxes according to a score function (as described below).
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionPartitionScoreSelectSampleEvaluate
Figure 2: Overview of the HOLLM algorithm: starting from initial data $\mathcal{D}$, it iteratively performs Partition, Score, Select, Sample (via LLM), and Evaluate steps to balance exploration and exploitation. For the partitioning here, we utilized a KD-Tree where each axis is split based on the mean values. Each rectangle represents a partition defined by the tree leaves. The red stars represent the new sampled points from the LLM.
3.2 Score: Synthesis of Exploitation, Geometry and Uncertainty
At every iteration $t$ the KD-tree yields a finite collection of leaves (partitions) $\{\mathcal{X}_{\ell,t}\}_{\ell=1}^{K_{t}}$. In order to decide where to spend our limited evaluation budget, we need to rank these leaves based on a scoring function (also called utility or acquisition function) that balances exploitation of good leaves and exploration of large regions that may hold good points and that are under-sampled.
(i) Exploitation via the empirical maximum. The exploitation term should optimistically reflect the best empirical evidence available for each region. Classical HOO algorithms use the sample mean as a low-variance proxy for the local reward (kleinberg08, ; wang08, ; bubeck11, ; valko13, ). In global optimization, however, the objective may be highly heteroscedastic, where one exceptionally good point inside an otherwise mediocre box can be more informative than the entire distribution. We therefore let our exploitation statistic be the largest improvement ever observed in a region $\mathcal{X}_{\ell,t}$:
$f_{\min}(t)=\min_{i\leq t}f(x_{i}),\quad Y_{i}=f(x_{i})-f_{\min}(t)+\varepsilon,\quad\mu_{\ell,t}=\max_{i\in I_{\ell,t}}Y_{i}.$ (1)
We subtract the current empirical minimum $f_{\min}(t)$ (since we are maximizing $f$) so the values become strictly non-negative and comparable across rounds  [^1] The additive constant $\varepsilon$ prevents zero scores during the startup phase.. Choosing a max rather than an average emphasizes regions that contain a good function value, a behavior also found in acquisition functions in Bayesian optimization (garnett_bayesoptbook_2023, ) and MCTS (wang20, ; yang21, ).
(ii) Geometric exploration through hypervolume. Let $[l_{\ell_{1}},u_{\ell_{1}}]\times\cdots\times[l_{\ell_{d}},u_{\ell_{d}}]$ be the axis-aligned hyperrectangle corresponding to leaf $\mathcal{X}_{\ell,t}$, where $l_{\ell_{d}}$ and $u_{\ell_{d}}$ are the low and upper axis values across dimension $d$ determined by the points in $\mathcal{X}_{\ell,t}=\{x_{i}\in\mathcal{X}:i\in I_{\ell,t}\}$. In order to assign a high exploration score to regions in the input space that are underexplored, we use the $d$-th root of the leaves’ Euclidean volume $\operatorname{Vol}(\mathcal{X}_{\ell,t})$: $V_{\ell,t}\;=\;\bigl{(}\,\prod_{j=1}^{d}(u_{\ell_{j}}-l_{\ell_{j}})\bigr{)}^{1/d}$, which is equivalent to the geometric mean of the side lengths of the hyperrectangle and is less sensitive to side lengths across single dimensions compared to the cell diameter. The $d$-th root scales $\operatorname{Vol}(\mathcal{X}_{\ell,t})$ so it has the same units as a length. Because axis-aligned boxes shrink anisotropically as the KD-tree refines, the $d$-th root removes the strong dependence on dimension and yields comparable numbers across $d$.
(iii) Statistical exploration via a UCB–V term. Even a tiny region may deserve further sampling if it contains a few samples with high variance. Let $\sigma_{\ell,t}^{2}$ be the empirical unbiased variance of the observed function values $\{Y_{i}\}_{i\in I_{\ell,t}}$ within region $\mathcal{X}_{\ell,t}$ at iteration $t$, and let $n_{\ell,t}=|I_{\ell,t}|$ be the number of samples in that cell. We adopt an exploration factor inspired by UCB-V (Upper Confidence Bound with Variance estimates) type algorithms (audibert08, ; audibert09, ; wang08, ; mukherjee18, ), and apply it to our dynamic KD-tree partitioning, reminiscing UCB-AIR for infinite-armed bandits where the number of arms increases at each iteration (wang08, ). More specifically, we score the region $\mathcal{X}_{\ell,t}$ with:
$\mathcal{E}_{\ell,t}=\sqrt{\frac{2\,\sigma_{\ell,t}^{2}\,\max\bigl{(}0,\ln(t/(K_{t}n_{\ell,t}))\bigr{)}}{n_{\ell,t}}}+\frac{c\cdot\max\bigl{(}0,\ln(t/(K_{t}n_{\ell,t}))\bigr{)}}{n_{\ell,t}}.$ (2)
[…]
Data: Initial data $\mathcal{D}$, budget $T$, batch size $b$, regions to sample from $M$, proposals per region $k$
1 while $t\leq T$ do
2       Update temperature $\alpha_{t}$ (and optionally maximum leaf size $m_{t}$)
       Partition space by building KD-tree on $\mathcal{D}$, obtaining $K_{t}$ leaves $\{\mathcal{\mathcal{X}}_{\ell,t}\}_{\ell=1}^{K_{t}}$
        // Partition
3
4      for each leaf $\mathcal{\mathcal{X}}_{\ell,t}$ do
5             Compute $\mu_{\ell,t}$ (Eq. 1), $V_{\ell,t}=\operatorname{Vol}(\mathcal{X}_{\ell,t})^{1/d}$ and $\mathcal{E}_{\ell,t}$ (Eq. 2)
             Normalize and compute total score $B_{\ell,t}=\bar{\mu}_{\ell,t}+\alpha_{t}\,\bigl{(}\beta_{1}\,\bar{V}_{\ell,t}+\beta_{2}\,\overline{\mathcal{E}}_{\ell,t}\bigr{)}$
              // Score
6
7       end for
      Select $M$ leaves by sampling with probabilities $p_{\ell,t}\propto B_{\ell,t}$
        // Select
8
[…]
        // Sample
9
10      Pick the top $b$ proposals by their LLM predicted scores
       Evaluate $f$ on them, add to $\mathcal{D}$, and set $t\leftarrow t+b$
        // Evaluate
11
12 end while
return best $(x,y)\in\mathcal{D}$
Algorithm 1 Hierarchical Optimization with LLMs (HOLLM)
Final composite score. All components must live on a shared numeric scale; otherwise, whichever component happens to have the largest dynamic range would dominate the others and nullify the intended trade–off. After each rebuild, we normalize the scores to $[0,1]$, preserving the intended relative weights even when the set of leaves changes drastically. The total score of each partition determined by the KD-tree partitioning is:
$B_{\ell,t}\;=\;\bar{\mu}_{\ell,t}+\alpha_{t}\,\bigl{(}\beta_{1}\,\bar{V}_{\ell,t}+\beta_{2}\,\overline{\mathcal{E}}_{\ell,t}\bigr{)},$ (3)
where $\bar{\mu}_{\ell,t},\,\bar{V}_{\ell,t},\overline{\mathcal{E}}_{\ell,t}$ are the min-max normalized scores, and $\beta_{1}$, $\beta_{2}$ are hyperparameters ($\beta_{1}+\beta_{2}=1$ by default) weighting the geometric versus statistical exploration. The $\alpha_{t}$ multiplier is a total exploration weight following an annealing schedule (cosine in our experiments). In the early phase ($\alpha_{t}\approx\alpha_{\max}$) the $B_{\ell,t}$ reduces to a near-uniform mixture of exploitation and the two exploratory terms. Assuming $\beta_{1}=\beta_{2}$ and non-drastically changing regions, as $t$ grows, the influence of $\bar{V}_{\ell,t}$ decays faster than that of $\overline{\mathcal{E}_{\ell,t}}$ because the latter itself shrinks with $n_{\ell}$. Hence, geometric exploration is front-loaded, while statistical calibration persists more throughout the optimization. When $t$ is close to $T$ the rule essentially becomes a greedy maximizer of $\bar{\mu}_{\ell,t}$, which is optimal once an $\varepsilon$-accurate maximizer has already been isolated. Thus, this composite score represents the classical trade-off: “go where I have seen something good, go where I have not looked at all, and go where my estimate is still uncertain”.
3.3 Select: Stochastic Selection of Partitions
Once the score $B_{\ell,t}$ (3) has been computed for every leaf, the algorithm must decide where to spend the next evaluation budget of size $b$. The Select step stochastically selects partitions by sampling from a categorical distribution over leaves. At round $t$, we draw without replacement a batch of $M$ distinct leaves, denoted as $\mathcal{B}_{t}$, from this categorical distribution where the sampling probability is: $p_{\ell,t}=B_{\ell,t}/\sum_{r=1}^{K_{t}}B_{r,t}$, where $\ell$ is the leaf index and $1\leq\ell\leq K_{t}$. Sampling stochastically instead of selecting the top-$M$ leaves means that sub-optimal leaves are sampled infinitely often (audibert08, ), potentially helping to mitigate premature convergence especially in highly non-convex and multimodal functions. Each leaf has always a positive probability due to the small constant $\epsilon>0$ we add to the exploitation term in Equation 1 and the min-max normalization in Equation 3. As $t$ grows, those exploratory components shrink and $B_{\ell,t}$ become increasingly peaked around the empirical best leaves, pushing $p_{\ell,t}$ toward a near-greedy regime. Moreover, a smooth annealing of $\alpha_{t}$ in Equation 3 avoids an abrupt "switch-to-greedy" policy, which may ignore late-appearing, high-value regions if they happen to be discovered just after the switch. Finally, sampling $M$ leaves without replacement diversifies evaluations by always sampling on distinct regions.
3.4 Sample: LLM-Guided Candidate Generation
After the Select step has identified a batch of leaves $\mathcal{B}_{t}=\{\mathcal{X}_{1,t},\dots,\mathcal{X}_{b,t}\}$, which also contain their corresponding hyperrectangular partition boundaries, HOLLM suggests new candidate points inside each chosen partition by prompting an LLM with the following logic: “Given the history of evaluations $\mathcal{D}_{t}$, propose $k$ new points that are likely to reveal high values of $f$ inside $\mathcal{X}_{i,t}$.” We construct a structured prompt (see Appendix D) containing: (i) points in $\mathcal{D}_{t}$ as in-context examples, (ii) the numeric partition bounds $(l_{i_{s}},u_{i_{s}})_{s=1}^{d}$ for cell $\mathcal{X}_{i,t}$, and (iii) task instructions to return new proposals and their estimated function values. Feeding this prompt to the LLM yields
$(\hat{\mathbf{x}}_{i},\hat{\mathbf{f}}_{i})\;=\;\textsc{LLM\_Generate}\!\bigl{(}\mathcal{D}_{t},(l_{i_{s}},u_{i_{s}})_{s=1}^{d},k\bigr{)},$
where $\hat{\mathbf{x}}_{i}=\{\hat{x}_{i,1},\dots,\hat{x}_{i,k}\}\subset\mathcal{X}_{i,t}$ and $\hat{\mathbf{f}}_{i}=(\hat{f}_{i,1},\dots,\hat{f}_{i,k})\in\mathbb{R}^{k}$ are, respectively, the LLM’s candidate locations and their predicted function values  [^4] We prompt the LLM to generate candidates and predict their performance with a single prompt or with two prompts, one for generation and one for prediction.. Across the $M$ selected leaves we thus obtain $k\cdot M$ suggestions. The parameter $k$ trades off the breadth of local exploration against prompt complexity and LLM inference cost. Finally, HOLLM keeps the globally best $b$ according to $\hat{f}$ and evaluates them on true function.
4 Empirical Evaluation
In this section, we evaluate HOLLM on a variety of search spaces and tasks. These span continuous synthetic functions and discrete search spaces for neural architecture search (NAS) (elsken2019neural, ; White2023NeuralAS, ) and hyperparameter optimization (Feurer2019, ; yu2020hyper, ).
Figure 3: Best function value across 100 iterations on the synthetic problems. HOLLM outperforms or matches the performance of baselines, especially on higher dimensional problems (e.g., Ackley).
Baselines. On these benchmarks, we compare against different state-of-the-art algorithms from Bayesian optimization (BO), such as multi-fidelity methods (CQR (salinas23a, )), Gaussian Process BO with Expected Improvement (GP-EI (snoek12, ; botorch, )), density estimator methods (TPE (NIPS2011_86e8f7ab, ) & BORE (tiao2021-bore, )), trust region BO (TuRBO eriksson19 ), evolutionary strategies (RE (real19, )) and random search (RS (bergstra12a, )). In all benchmarks, we also compare to the global LLM-based optimizer baseline (see Algorithm 2 in appendix) that uses the exact same prompt structure as HOLLM (we provide the prompt templates in Appendix D), with the only difference being the region boundaries.
Setup. Starting from $n_{0}=5$ initial random evaluations, we run each method 3 times for a total of $T=100$ iterations with different random seeds and report their mean and standard error. We use their implementation in SyneTune (salinas22a, ), except TuRBO, for which we use the official BoTorch code from the authors (botorch, ). If not stated otherwise, for HOLLM we always decay the $\alpha_{t}$ exploration coefficient from $1.0$ to $0.01$ using a cosine annealing schedule (loshchilov2017sgdr, ), a batch size $b=4$, $M=5$ selected partitions, $k=5$ proposals per selected partition, and a fixed maximum leaf size $m_{t}=m_{0}=\lceil d/2\rceil$. In Appendix C.1, we provide ablations on these hyperparameter choices in our algorithm. We use Gemini-1.5-Flash as the LLM in Sample due to its fast inference speed, low cost, and large context window. Importantly, the LLM is provided with only minimal task information: the input dimensionality, variable names whenever applicable (e.g., hyperparameter names), partition boundaries, and in-context examples. No task-specific descriptions or dataset statistics are included. While prior work (liu2024large, ) shows that performance can improve by enriching prompts with such information, we avoid this to prevent potential contamination and reported performance on overly engineered prompts. We provide the full experimental details in Appendix B.
Figure 4: Illustrative example of HOLLM optimizing a 1D multimodal problem. The rectangles represent the space partitions (top figure) and are highlighted in orange whenever they are selected based on their respective probabilities (bottom figure). We used a batch size of 3. All new points (red stars) are LLM suggestions. Notice how partitions become fine-grained around the global maximum.
4.1 Synthetic Functions
We benchmark on six synthetic functions of varying nature and dimensionality: Hartmann-3D and Hartmann-6D (smooth but sharply multimodal), Rosenbrock-8D (unimodal with a narrow valley and ill-conditioning), Rastrigin-10D (regular multimodality with $10^{10}$ local minima), Lévy-10D (plateaus, cliffs, and funnels), and Ackley-20D (flat regions and a single sharp global minimum at the origin). These functions pose challenges ranging from fine local search to broad exploration. See Table 1 in Appendix B.2 for more details on these functions. Results presented in Figure 3 show that HOLLM consistently outperforms the global LLM baseline, especially on the multimodal functions, also exhibiting less variance between runs. It also matches or surpasses all other baselines. Most notably, on Ackley-20D with input range $[-32.768,32.768]^{20}$, HOLLM locates the global maximum in just 50 iterations, while baselines struggle to improve beyond random search.
Figure 5: Hyperparameter optimization on 4 datasets from the FCNet search space. All baselines from Synetune are evaluated asynchronously using 4 workers.
Visualizing the Optimization Process. In Figure 4, we show a visualization of HOLLM’s mechanics on a 1D multimodal function. The rectangles represent the KD-tree space partitions and they are highlighted in orange whenever they get selected. We can see that during the first iterations the partitions are larger and HOLLM is more exploratory, also confirmed by the regions’ respective probabilities (bottom bar plot). Later on, as the regions become smaller, high modes are identified and by the end the score probability mass concentrates more around the global maximum. We provide similar visualizations for Lévy-1D and Rosenbrock-1D in Appendix C.
4.2 Hyperparameter Optimization
We assess the effectiveness of HOLLM on hyperparameter optimization by optimizing the 9D categorical space from FCNet klein2019 , where the task is to minimize the validation MSE of a fully connected network on 4 distinct datasets: PROTEIN, NAVAL, PARKINSONS and SLICE. See Appendix B.3 for more details on this search space. Results shown in Figure 5 demonstrate that our method outperforms or is on par with methods such as BORE and CQR, which typically are the off-the-shelf best choices on these benchmarks. Compared to the global LLM baseline, we can clearly see improvements on all datasets except on Parkinson, where the LLM seems to benefit more by sampling globally and reaches a low MSE after only 20 iterations. This may be due to potential outliers in the data that may impact HOLLM’s performance.
4.3 Neural Architecture Search
Figure 6: Results on the NAS-Bench-201 6 dimensional discrete function. We plot the negative regret vs. the number of iterations. We run each method 6 times and report the mean and standard error.
Neural Architecture Search (NAS) (White2023NeuralAS, ), like hyperparameter optimization, aims to identify the best-performing neural network architecture for a given dataset by maximizing validation accuracy. We use the NAS-Bench-201 benchmark (nb201, ), which provides precomputed validation accuracies for all architectures on CIFAR-10, CIFAR-100, and Downsampled ImageNet $16\times 16$ (imgnet, ). The search space is 6D, with each dimension representing a discrete choice among 5 possible layer operations. See Appendix B.4 for full details. We use a continuous representation $[0,1]^{6}$ of the input space and discretize it to evaluate the true function. As seen in Figure 6, HOLLM always outperforms the LLM baseline that samples globally and is on par with BORE and CQR. The global LLM seems to get stuck in local minima, therefore leading to stagnating performance from early on.
Figure 7: Ablations on HOLLM’s components.
Ablations. To assess the impact of key design choices in HOLLM, we perform the following ablations: (1) we modify the score function in Equation 3 by isolating either the exploitation or exploration term; (2) we replace the variance-aware UCB-V bonus in Equation 2 with the simpler UCB1 (auer02, ); (3) we substitute the categorical distribution used in line 8 of Algorithm 1 with a uniform distribution; and (4) we replace the LLM-based sampler in Sample with non-LLM baselines such as uniform sampling and samples from a local Gaussian Process fitted to each partition. Results in Figure 7 show that the choice of candidate sampler in Sample has the most significant effect on regret.
5 Conclusion, Limitations and Societal Impact
We propose HOLLM, a novel LLM-based global optimization method for expensive blackbox functions, that combines adaptive KD-tree partitioning, a bandit-inspired score function, and LLM capabilities for generating new candidate points on locally promising regions. HOLLM excels especially on multimodal functions with many local optima that pose a risk for premature convergence, hyperparameter optimization and neural architecture search, consistently outperforming LLMs with a global sampling policy and other non-LLM state-of-the-art methods.
Limitations. While HOLLM combines nonparametric bandit methods with LLM-based sampling and shows strong empirical performance, it has several limitations. First, the approach currently lacks formal theoretical guarantees, particularly regarding dynamic partitioning and regret bounds, which we leave for future work. Second, its effectiveness depends heavily on the quality of LLM-generated proposals; biased or miscalibrated models can misguide the search or waste evaluations. Third, the inference and monetary cost of LLMs, especially proprietary ones, can limit scalability in high-dimensional settings. Finally, although default parameter values perform well in our experiments, real-world deployment may require tuning them to avoid premature convergence or excessive exploration.
Impact. The use of LLMs in global optimization has societal implications. HOLLM has the potential to accelerate progress in areas such as drug discovery, materials design, and energy systems by reducing experimental costs and enabling personalized solutions. On the other hand, reliance on LLMs trained on biased data risks perpetuating social injustices when guiding sensitive decisions (e.g., hiring). Additionally, repeated LLM queries incur considerable energy costs, and the opacity of LLM-driven decisions may limit transparency and reproducibility. Therefore, responsible deployment requires bias assessment, usage controls, and transparency in both computational and ethical impacts.
Acknowledgments
Robert Bosch GmbH is acknowledged for financial support. Fabio Ferreira and Frank Hutter acknowledge funding by the European Union (via ERC Consolidator Grant DeepLearning 2.0, grant no. 101045765). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. Aaron Klein acknowledges the financial support by the Federal Ministry of Education and Research of Germany and by Sächsische Staatsministerium für Wissenschaft, Kultur und Tourismus in the programme Center of Excellence for AI-research „Center for Scalable Data Analytics and Artificial Intelligence Dresden/Leipzig", project identification number: ScaDS.AI. Frank Hutter acknowledges the financial support of the Hector Foundation. We also thank Google Cloud for their free trial program, that enabled us to use the Google Gemini models throughout this project.
Appendix A Algorithm Pseudocodes
In this section, we present detailed pseudocode for the HOLLM algorithm in Algorithm 3, which complements Algorithm 1. In the algorithm, we omit the subscript $t$ for easier readability. Additionally, Algorithm 2 describes the LLM-based global optimization baseline method used throughout our experiments. To ensure fair comparison, we configure the baseline to propose $k\cdot b$ points per iteration (line 2 of Algorithm 2), matching the total number of proposals generated by HOLLM across all subregions, i.e., $k$ proposals in each of the $M$ subregions.
Data: Initialize $\mathcal{D}$ with $n_{0}$ points, budget $T$, batch size $b$, proposals $k\cdot M$
1
2for $t=n_{0},\dots,T-1$ do
3       Propose $k\cdot M$ points with LLM_Generate$(\mathcal{X},\mathcal{D},k\cdot M)$
4       Evaluate the top $b$ and add to $\mathcal{D}$
5
return best $(x,y)\in\mathcal{D}$
Algorithm 2 Global-LLM baseline
Data: Initial data $\mathcal{D}=\{(x_{i},f(x_{i}))\}_{i=1}^{n_{0}}$, batch size $b$, regions to sample from $M$, proposal count per leaf $k$, dimension $d$, initial leaf size $m_{0}$, adaptive growth rate $\lambda$, total evaluations $T$, exploration weights $\beta_{1}$, $\beta_{2}$, annealing $\alpha_{min}$, $\alpha_{max}$
1
$t\leftarrow n_{0}$
  // global evaluation counter
2 while $t<T$ do
       (optional) $m_{\text{leaf}}\leftarrow m_{0}+\lceil\lambda\log(1+t)\rceil$
        // adaptive leaf size
3
4      Build KD-tree on $\mathcal{D}$ with leafsize $m_{\text{leaf}}$, obtaining $K_{t}$ leaves $\{\mathcal{\mathcal{X}}_{\ell}\}_{\ell=1}^{K_{t}}$
5
      $\alpha\leftarrow\alpha_{min}+\frac{1}{2}(\alpha_{max}-\alpha_{min})(1+\cos(\pi t/T))$
        // cosine annealing schedule
6
7      $f_{\min}\leftarrow\min_{i=1}^{t}f(x_{i})$
       $Y_{+}\leftarrow\{f(x_{i})-f_{\min}+\epsilon\}_{i=1}^{t}$
        // positive transformed values
8
9      for $l=1$ to $K_{t}$ do
             $n_{\ell}\leftarrow|\mathcal{X}_{\ell}|$
              // number of points in this leaf
10
11            $(l_{\ell},u_{\ell})\leftarrow$ bounds of cell $\mathcal{\mathcal{X}}_{\ell}$
             $V_{\ell}\leftarrow\prod_{j=1}^{d}(u_{\ell j}-l_{\ell j})^{1/d}$
              // normalized volume
12
13            if $n_{\ell}>0$ then
                   $\mu_{\ell}\leftarrow\max_{i\in I_{\ell}}Y_{+}[i]$
                    // best value in cell
14                   if $n_{\ell}>1$ then
                         $\sigma_{\ell}^{2}\leftarrow\frac{1}{n_{\ell}-1}\sum_{i\in I_{\ell}}(Y_{+}[i]-\bar{Y}_{\ell})^{2}$
                          // variance in cell
15
16                  else
                         $\sigma_{\ell}^{2}\leftarrow 0.01$
                          // default variance for single-point cells
17
18
19                  $\log\_term\leftarrow\max(0,\log(t/(K_{t}\cdot n_{\ell})))$
20                   $\mathcal{E}_{\ell}\leftarrow(\sqrt{2\sigma_{\ell}^{2}\cdot\log\_term/n_{\ell}}+\log\_term/n_{\ell})$
21
22            else
23                   $\mu_{\ell}\leftarrow 0$
                   $expl_{\ell}\leftarrow 1$
                    // high exploration for empty cells
24
25
26            $\bar{\mu},\bar{V_{\ell}},\bar{\mathcal{E}_{\ell}}\leftarrow$ min-max normalize $\mu_{\ell}$, $V_{\ell}$, $\mathcal{E}_{\ell}$ across all cells
27
            $B_{\ell}\leftarrow\bar{\mu}+\alpha\cdot(\beta_{1}\cdot\bar{V_{\ell}}+\beta_{2}\cdot\bar{\mathcal{E}_{\ell}})$
              // composite score
28
29
      $p_{\ell}=B_{\ell}/\sum_{r=1}^{K_{t}}B_{r}$
        // normalize score across cells
30       Sample $M$ cells $\{\mathcal{\mathcal{X}}_{i_{j}}\}_{j=1}^{M}\sim\mathrm{Categorical}\bigl{(}\{p_{\ell}\}\bigr{)}$
31
32      $\hat{X}\leftarrow\emptyset,\ \hat{F}\leftarrow\emptyset$
33       for $j=1$ to $M$ do
34             $(\hat{\mathbf{x}}_{j},\hat{\mathbf{f}}_{j})\leftarrow\textsc{LLM\_Generate}(\mathcal{D},(l_{i_{j}},u_{i_{j}}),k)$
35             Append $\hat{\mathbf{x}}_{j}$ to $\hat{X}$; append $\hat{\mathbf{f}}_{j}$ to $\hat{F}$
36
37
      $\pi\leftarrow\text{argsort}(\hat{F})$
        // indices of sorted values (descending)
38       $X^{\text{new}}\leftarrow$ top $b$ points from $\hat{X}$ using indices $\pi$
39
40      for each $x\in X^{\text{new}}$ do
41             Evaluate $y=f(x)$
42             $\mathcal{D}\leftarrow\mathcal{D}\cup\{(x,y)\}$
43             $t\leftarrow t+1$
44
45
return best point $(x^{*},f(x^{*}))$ where $x^{*}=\arg\max_{x\in\mathcal{D}}f(x)$
Algorithm 3 Hierarchical Optimization with LLMs (HOLLM) – Detailed
Appendix B Details on Tasks, Baselines and Experimental Setup
B.1 Baselines
We compare HOLLM to the following baselines:
  • Random Search (RS) [8] serves as a simple baseline that uniformly samples configurations from the search space without any learning or adaptation.
  • Regularized Evolution (RE) [42] is an evolutionary algorithm that maintains a population of candidate solutions and evolves them through mutation operations. The method regularizes the population by removing the oldest individuals, preventing stagnation and maintaining diversity.
  • Conformalized Quantile Regression (CQR) [45] uses gradient boosted trees to predict performance quantiles and provides prediction intervals with statistical guarantees through conformal prediction techniques.
  • Tree-structured Parzen Estimator (TPE) [7] is a sequential model-based optimization method that models the distribution of good and bad configurations separately. It builds two probability densities: $\ell(x)$ for configurations with performance better than a threshold, and $g(x)$ for those below the threshold. New candidates are selected by maximizing the ratio $\ell(x)/g(x)$.
  • Bayesian Optimization by Density-Ratio Estimation (BORE) [52] reformulates Bayesian optimization as a binary classification problem. It trains a probabilistic classifier to distinguish between high-performing and low-performing configurations, then uses the predicted class probabilities to construct an acquisition function equivalent to expected improvement.
  • Gaussian Process BO with Expected Improvement (GP-EI) [50, 6] employs a Gaussian Process as a surrogate model to capture the objective function’s behavior and uncertainty. It uses the Expected Improvement acquisition function, implemented via BoTorch, to balance exploration and exploitation when selecting new evaluation points.
  • Trust Region Bayesian Optimization (TuRBO) [18] addresses the curse of dimensionality in high-dimensional optimization by maintaining multiple local trust regions. Each region uses an independent Gaussian Process and adapts its size based on optimization progress, allowing the method to scale effectively to high-dimensional problems.
B.2 Synthetic Benchmarks
In Section 4, we initially evaluated HOLLM and the baselines on 6 synthetic deterministic function with varying dimensionality and nature. In Table 1, we provide details on each of them.
Table 1: List of synthetic optimization functions and their main characteristics.
Function (dim.) Landscape & key traits Global boundary Global optimum
Hartmann 3D Smooth, strongly multimodal surface generated by four weighted Gaussians inside the unit cube; narrow, steep basins punish local search. $(x_{1},x_{2},x_{3})\in[0,1]^{3}$ $f_{\min}\!\approx\!-3.86278$
Hartmann 6D Six Gaussians in $[0,1]^{6}$ create an even denser constellation of deceptive wells; still smooth but mildly ill-conditioned, and the search space grows exponentially. $(x_{1},x_{2},\dots,x_{6})\in[0,1]^{6}$ $f_{\min}\!\approx\!-3.32237$
Rosenbrock 8D Classic curved “banana” valley stretched to eight variables; unimodal yet highly ill-conditioned, requiring precise valley-tracking; non-separable. $(x_{1},x_{2},\dots,x_{8})\in[-2.048,2.048]^{8}$ $f_{\min}=0$
Rastrigin 10D Quadratic core overlaid with cosine ripples forms a perfectly regular grid of $10^{10}$ local minima; separable but brutally multimodal, exposing algorithms prone to premature convergence. $(x_{1},x_{2},\dots,x_{10})\in[-5.12,5.12]^{10}$ $f_{\min}=0$
Lévy 10D Sine perturbations on a quadratic backbone yield wide plateaus, sudden cliffs, and deep funnels—rugged and non-separable, stressing step-size control. $(x_{1},x_{2},\dots,x_{10})\in[-10,10]^{10}$ $f_{\min}=0$
Ackley 20D Exponential of radius plus averaged cosines: vast flat outer region, encircling ridge, and a single sharp basin at the origin; tests exploration versus exploitation in very high dimension. $(x_{1},x_{2},\dots,x_{20})\in[-32.768,32.768]^{20}$ $f_{\min}=0$
B.3 Hyperparameter Optimization Benchmarks
For our hyperparameter optimization experiments, we evaluate on four tasks from the FCNet benchmark [28]: PROTEIN, NAVAL, PARKINSONS, and SLICE. The FCNet benchmark provides a tabulated hyperparameter optimization setting where fully connected neural networks are trained on each dataset with different hyperparameter configurations. The search space consists of 9 categorical hyperparameters (network architecture and training parameters), yielding 62,208 possible configurations with pre-computed validation accuracies. To enable KD-tree partitioning on the categorical search space, we apply ordinal encoding to convert categorical variables into numerical split indices. Below we describe the four regression datasets used as the underlying machine learning tasks:
  • PROTEIN is a regression dataset containing physicochemical properties of protein tertiary structures. The task involves predicting protein properties from 9 molecular descriptors across 45,730 protein samples.
  • PARKINSONS contains biomedical voice measurements from 42 individuals with early-stage Parkinson’s disease participating in a six-month telemonitoring trial. The regression target is the progression of Parkinson’s symptoms, with 5,875 samples and 19 acoustic features.
  • NAVAL consists of simulated sensor data from a naval frigate’s propulsion system, including gas turbine, propeller, gearbox, and control systems. The regression task predicts component degradation levels using 11,934 samples with 16 operational features.
  • SLICE involves predicting the relative axial location of CT scan slices within the human body. The dataset contains 384 features extracted from 53,500 CT images, describing bone structures, air inclusions, and anatomical positioning.
Table 2: Search space of the FCNet benchmark. The left column lists the hyperparameter names of the neural network that need to be tuned, whilst the right column the possible categorical choices for each hyperparameter.
Hyperparameter Categorical Configuration Space
Initial LR {0.0005, 0.001, 0.005, 0.01, 0.05, 0.1}
Batch Size {8, 16, 32, 64}
LR Schedule {cosine, fix}
Activation (Layer 1) {relu, tanh}
Activation (Layer 2) {relu, tanh}
Layer 1 Size {16, 32, 64, 128, 256, 512}
Layer 2 Size {16, 32, 64, 128, 256, 512}
Dropout (Layer 1) {0.0, 0.3, 0.6}
Dropout (Layer 2) {0.0, 0.3, 0.6}
B.4 Neural Architecture Search Benchmarks
For neural architecture search (NAS), we utilize the NAS-Bench-201 [16] tabular benchmark, which provides a comprehensive evaluation suite for architecture optimization. The search space consists of selecting optimal CNN operations for each of the 6 edges in a predefined cell-based computational graph. Each edge can be assigned one of 5 categorical operations: avg_pool_3x3 (average pooling), nor_conv_3x3 (normal 3×3 convolution), skip_connect (identity connection), nor_conv_1x1 (normal 1×1 convolution), and none (no operation). This yields a total search space of $6255^{6}=15{,}$ 625 possible architectures. NAS-Bench-201 provides precomputed validation accuracies for all architectures across three image classification datasets: CIFAR-10, CIFAR-100, and ImageNet16-120 (a 16×16 downsampled version of ImageNet with 120 classes). This tabulated format enables efficient benchmarking by eliminating the computational overhead of training each architecture from scratch.
Appendix C Additional Experiments
In this section, we provide additional experiments and ablations, complementing the ones conducted throughout Section 4 of the main paper.
(a) 1D Levy problem.
(b) 1D Rosenbrock problem.
Figure 8: Illustrative examples of optimizing 1D functions. The rectangles represent the input space partitions (top row). They are highlighted in orange whenever selected based on their respective probabilities (bottom row). Both experiments used 5 initial points and a batch size of 4. All new points (red stars) are suggestions from the LLM.
C.1 Ablations
To assess the robustness of our method and understand the influence of key hyperparameters on performance, we conducted a comprehensive ablation study. We employ the 10D Levy test function and examine 3 hyperparameters that directly impact the exploration-exploitation balance and efficacy of our approach: (i) maximum leaf capacity $m_{\text{leaf}}=m_{0}+\lceil\lambda\log(1+t)\rceil$ ($\lambda=0$), which controls the granularity of space partitioning; (ii) candidate sampling rate $k$ (proposals generated per selected region), which determines the diversity of proposals within each selected region; and (iii) region selection parameter $M$ (partitions selected per iteration), which governs the number of promising subregions explored simultaneously per iteration. The default hyperparameter configuration also used throughout the experiments in the main paper is: exploration parameter bounds $\alpha_{\text{max}}=1.0$ and $\alpha_{\text{min}}=0.01$, initial random sampling phase of $n_{0}=5$ evaluations, batch size $b=4$ (points evaluated per iteration), $k=5$, $M=5$, and maximum leaf capacity $m_{0}=d/2$, where $d$ denotes problem dimensionality. We run each setting with 5 independent random seeds and report the mean performance $\pm$ standard error in Figure 9.
Figure 9: Impact of key hyperparameters on optimization performance for the 10D Levy function. Left: Ablation on leaf size ($m_{0}$) demonstrates that coarse partitioning ($m_{0}=1d$) yields the poorest performance, while finer partitioning enables superior exploitation of promising regions through more granular space decomposition. Middle: Varying the number of candidates per selected region ($k$) reveals optimal performance at intermediate values, where undersampling ($k=1$) significantly degrades performance by limiting exploitation of high-potential regions, and oversampling ($k=10$) slows convergence due to inefficient allocation of evaluation budget. Right: The number of partitions selected per trial ($M$) governs exploration breadth, where single-region focus ($M=1$) impedes convergence through insufficient exploration, while moderate values ($M\in\{3,5,7\}$) accelerate optimization by enabling simultaneous exploration of multiple promising regions.
Impact of Leaf Size ($m_{0}$). The leaf size parameter $m_{0}$ defines the maximum number of data points within a single leaf of the partitioning tree, directly controlling the granularity of search space decomposition. Our analysis across different values of $m_{0}$ as a factor of problem dimensionality $d$ reveals a clear trade-off between partition resolution and statistical reliability (Figure 9, left). Coarse partitioning with $m_{0}=d$ yields suboptimal performance due to overly broad regions that group diverse areas of the search space, diminishing the method’s ability to precisely isolate promising subregions. Conversely, extreme fine partitioning with $m_{0}=1$ also degrades performance because singleton regions provide insufficient statistical information and the variance component becomes a small constant across all regions, eliminating valuable uncertainty estimates necessary to guide exploration. We observe the best performance at $m_{0}=d/4$, which strikes an effective balance by enabling detailed space partitioning while maintaining sufficient data density within each region to compute meaningful variance estimates for the exploration term.
Impact of Number of Candidates per Region ($k$). We investigated the effect of varying the number of candidate points $k$ sampled from each selected region, testing values $k\in\{1,3,5,7,10\}$. Results shown in Figure 9, middle) reveal a clear trade-off between under- and over-sampling within regions. Setting $k=1$ leads to significant performance degradation as the method fails to adequately exploit promising regions by drawing only a single sample per region. Conversely, $k=10$ results in worse performance during initial iterations compared to intermediate $k$ values, which we attribute to increased risk of oversampling sub-optimal regions in the beginning. While the method can recover from this scenario as oversampled sub-optimal regions receive lower scores in subsequent iterations, the initial performance penalty demonstrates that excessive sampling can be counterproductive. These findings showcase the importance of balanced exploitation within selected regions: sufficient sampling to capitalize on promising areas without overcommitting computational budget to potentially sub-optimal regions.
Impact of Number of Partitions Selected per Trial ($M$). We examined how the number of partitions $M\in\{1,3,5,7\}$ selected per trial affects optimization performance. As seen in Figure 9, right, setting $M=1$ notably hinders performance, particularly during initial iterations, as HOLLM severely restricts exploration breadth by focusing all sampling efforts on a single region per iteration. When the initially chosen region lacks global promise, progress becomes slow, though the method can eventually exploit good regions once identified, leading to convergence that is typically slower than broader exploration strategies. Conversely, increasing $M$ to moderate or high values (3, 5, or 7) generally improves initial exploration by enabling simultaneous consideration of multiple diverse regions. This expansion of the candidate selection pool allows the algorithm to benefit from a larger, more diverse set of proposals per trial, improving performance up to a saturation point. The results demonstrate that balanced multi-region exploration through appropriate $M$ values provides superior performance compared to overly focused single-region strategies, highlighting the importance of maintaining exploration breadth while preserving the ability to exploit promising areas effectively.
C.1.1 Impact of exploration parameter $\alpha_{max}$
Figure 10: Performance comparison of exploration parameter settings ($\alpha_{\text{max}}\in\{0.2,0.5,0.7,1.0\}$) across the four FCNet benchmark tasks. Each curve represents the mean objective function value over 5 independent runs, with shaded regions denoting standard error. The results illustrate that optimal $\alpha_{\text{max}}$ selection exhibits task-dependent behavior, reflecting the varying landscape characteristics and exploration requirements across different hyperparameter optimization problems.
We evaluated the effect of different exploration settings on the FCNet benchmarks across four tasks: PROTEIN, NAVAL, PARKINSONS, and SLICE. Results in Figure 10 show that the impact of the exploration parameter $\alpha_{\text{max}}$ exhibits task-dependent variation, with optimal settings determined by the underlying problem structure. Higher values of $\alpha_{\text{max}}$ bias the search toward less-explored regions, proving beneficial for highly multimodal or non-convex landscapes where diverse exploration is crucial for escaping local optima. Conversely, lower $\alpha_{\text{max}}$ values reduce exploration of new regions and concentrate search efforts on exploitation, which may be more appropriate for smoother or convex solution spaces where intensive local search around promising areas yields better returns. These findings suggest that prior knowledge of the task landscape characteristics, such as modality, convexity, and noise structure, can effectively guide the selection of $\alpha_{\text{max}}$ to match the exploration-exploitation balance to the problem’s inherent difficulty and structure.
C.1.2 Effect of Hyperparameters on Computational Cost
The computational complexity of our method scales directly with the total number of candidate points generated per iteration, calculated as $k\cdot M$. When employing computationally expensive models such as large language models accessed via API calls or hosted locally, increases in either $M$ (which amplifies the number of inference calls per iteration) or $k$ (which extends the number of output tokens generated per call) result in proportionally higher inference times and associated costs per optimization step. This creates a fundamental trade-off between optimization performance and computational efficiency that practitioners must carefully consider based on their specific resource constraints and performance requirements. Our default configuration of $M=5$ and $k=5$, yielding 25 candidate evaluations per iteration, represents a calibrated compromise that balances exploration capability with computational practicality across the diverse benchmarks in our experimental evaluation.
Appendix D Prompts
We design structured prompts to LLMs for both candidate generation and evaluation prediction within our optimization framework. Our prompting strategy consists of two complementary components: a candidate generation prompt (3) that produces new solutions based on historical observations, and an evaluation prediction prompt (4) that estimates the quality of proposed candidates, or both in a single prompt. Each prompt follows a systematic structure:
  1. 1.
    Task specification: We provide a description of the optimization objective and establish the problem context, including the nature of the search space and optimization goals.
  2. 2.
    Dynamic constraints: We define the feasible region constraints (boundaries) derived from the current KD tree partitioning. These constraints are automatically computed based on the selected leaf nodes and translated into natural language descriptions that specify valid input ranges alongside example-evaluation pairs.
  3. 3.
    In-context examples: We supply the model with historical observations consisting of previously evaluated points and their corresponding objective function values. These examples serve as demonstrations to guide the model’s understanding of the optimization landscape and desired output format.
  4. 4.
    Task-specific instructions: We provide explicit directives tailored to each prompt’s purpose. The candidate generation prompt instructs the model to propose new configurations, while the evaluation prediction prompt directs it to estimate performance for given candidates. Both prompts enforce structured JSON output formatting for automated parsing of model responses.
Throughout our prompts, we employ placeholder variables denoted with \$ symbols to represent task-specific information that is dynamically populated during execution. Comprehensive examples and descriptions for each placeholder are provided in Table D. For the NAS-Bench-201 experiments, we adopt a streamlined approach using a unified prompt structure (5) that simultaneously elicits both candidate proposals and their predicted evaluations in a single model query, reducing the computational overhead of separate generation and prediction phases.
Table 3: Description of placeholders for candidate proposal and prediction prompts.
Placeholder Description Example of Replaced Text
\$metrics The performance metrics for the specific task. F1 (lower is better)
[…]
  lr: range(float([0.0, 0.9])),
  activation: choice(["relu", "tanh"]),
  num_layer: range(int([1, 20]))
[…]
   {lr: 0,4, activation: "relu", num_layer: 8,...}
   F1: 5.65
   {lr: 0.03, activation: "tanh", num_layer: 8,...}
   F1: 3.23
[…]
  lr: ?,
  activation: ?,
  num_layer: ?
[…]
   1: {lr: 0,4, activation: "relu", num_layer: 8,...}
   2: {lr: 0.03, activation: "tanh", num_layer: 8,...}
[…]
  {F1: ? }
Suggest 100 random samples for 8 dimensions within the specified bounding box with a maximum of 3 decimal places.
Bounding Box:
x1_min: 0, x1_max: 1
x2_min: 0, x2_max: 1
x3_min: 0, x3_max: 1
x4_min: 0, x4_max: 1
x5_min: 0, x5_max: 1
x6_min: 0, x6_max: 1
x7_min: 0, x7_max: 1
x8_min: 0, x8_max: 1
Return the suggestions in the following JSON format exactly, without any additional text:
[{"x1": float, "x2": float, "x3": float, "x4": float, "x5": float, "x6": float, "x7": float, "x8": float}]
Listing 1: Prompt for LLMs simulating 100 8-D uniform random samples.
Suggest 80 sample points in 2 dimensions within the specified bounding box.
Bounding Box:
x1_min: 0, x1_max: 1
x2_min: 0, x2_max: 1
such that they are clustered around the points given below.
Points:
Point 1: x1: 0.25, x2: 0.25
Point 2: x1: 0.75, x2: 0.75
Return the suggestions in the following JSON format exactly, without any additional text:
[{"x1": float, "x2": float}]
Listing 2: Prompt for LLMs sampling 80 points around the minima.
# Optimization task
## Problem Description
You are tasked with solving a optimization problem that requires finding optimal solutions.
- **Evaluation**: Configurations are measured by \$metrics
## Constraints
The allowable ranges for the hyperparameters are:
\$region_constraints
## Previously Evaluated Architectures
Below are examples of architectures that have been evaluated, showing their operations and performance metrics:
\$Region_ICL_examples
## Your Task
Generate \$target_number_of_candidates new architecture configurations that:
1. Are likely to achieve lower \$metrics than the examples
2. Are different from all previously evaluated architectures
3. Satisfy all the specified constraints: \$region_constraints
## Output Format
Each configuration has to follow this format:
\$candidate_sampler_response_format
Provide your response in a JSON list containing each proposed configuration.
Return only the required JSON list output without additional text.
Listing 3: Prompt template used for candidate points generation in the LLM_Generate function.
# Configuration Performance Prediction
## Problem Description
You are tasked with predicting the performance of configurations.
- **Evaluation Metric**: \$metrics (to be predicted)
- **Constraint**: The allowable ranges for the hyperparameter are: \$Region_ICL_examples
## Reference configurations with Known Performance
Below are examples of configurations that have been evaluated, showing their operations and performance metrics:
## Candidate configurations to Evaluate
You must predict performance for these new configurations:
\$target_architectures
## Your Task
1. Predict the \$metrics value for each candidate configuration
2. Base your predictions on patterns in the reference examples
## Output Format
Each evaluation has to follow this format:
\$surrogate_model_response_format
Provide your response in a JSON list containing each proposed evaluation.
Return only the required JSON list output without additional text.
Listing 4: Prompt template used for performance prediction in the LLM_Generate function.
Suggest 8 new candidate point(s) for maximizing a blackbox function in a 6-dimensional search space.
Below are some examples of previously evaluated points with their corresponding function values:
"x1": 0.034,
"x2": 0.287,
"x3": 0.773,
"x4": 0.175,
"x5": 0.755,
"x6": 0.608,
"value": -37.093
"x1": 0.199,
"x2": 0.433,
"x3": 0.405,
"x4": 0.779,
"x5": 0.186,
"x6": 0.594,
"value": -37.84
"x1": 0.447,
"x2": 0.342,
"x3": 0.97,
"x4": 0.087,
"x5": 0.115,
"x6": 0.533,
"value": -44.52
"x1": 0.949,
"x2": 0.127,
"x3": 0.659,
"x4": 0.546,
"x5": 0.049,
"x6": 0.265,
"value": -33.067
The search space is defined by the following bounding boxes:
x1_min: 0.492, x1_max: 1.000
x2_min: 0.000, x2_max: 1.000
x3_min: 0.000, x3_max: 1.000
x4_min: 0.000, x4_max: 1.000
x5_min: 0.000, x5_max: 1.000
x6_min: 0.000, x6_max: 1.000
Based on the examples above, suggest candidate points that balance exploration (sampling new regions) with exploitation (focusing on promising areas where function values are good). Each candidate point must lie within the specified bounding boxes. In addition, predict an estimated function value for each candidate.
Return the suggestions in the following JSON format exactly, without any additional text:
[{"x1": float, "x2": float, "x3": float, "x4": float, "x5": float, "x6": float, "value": float}]
Listing 5: Prompt example used for simultaneous candidate generation and performance prediction in the LLM_Generate function for NAS-Bench-201.